OpenStudy (anonymous):

I have the answer cansomeone please tell me what they got... Find the derivative of k(x)=tan(e^x)

OpenStudy (ash2326):

\[k(x)=\tan (e^x)\] \[k'(x)=\sec^2 (e^x)\times \frac{d}{dx} (e^x)\] Did you do this?

OpenStudy (anonymous):

yes and I got for an answer\[e^xcos^2(e^x)\]

OpenStudy (ash2326):

it should be \(\sec^2 (e^x)\)

OpenStudy (anonymous):

could it be cos^(e^x) I don't have sec as a choice

OpenStudy (ash2326):

cos^2 e^x could be in the denominator \[\frac{e^x}{\cos^2{e^x}}\]

OpenStudy (anonymous):

ok yes I have that for one of my choices. So the answer that you just posted is another way in saying the answer correct.

OpenStudy (ash2326):

yes since \[\sec x =\frac{1}{\cos x}\]

OpenStudy (anonymous):

ok gotcha thanks

OpenStudy (ash2326):

Welcome :)